OpenStudy (anonymous):

The 5th term in a geometric sequence is 120. The 7th term is 30. What are the possible values of the 6th term of the sequence?

OpenStudy (johnweldon1993):

Well if we were going the other way....we could see that 30 * 2 = 60 60 * 2 = 120 so that would be a way to approach it and see that a possible answer for the 6th term could be 60 and the common ratio would be 1/2 :)

OpenStudy (anonymous):

\[A.60 B.\pm60C.75D.\pm75\] are my choices

OpenStudy (whpalmer4):

You're trying to find 1 geometric mean between 120 and 30. 120:x = x:30 or \[\frac{120}{x}= \frac{x}{30}\]\[x^2=3600\]\[x=60\]

OpenStudy (radar):

120r^2= 130 \[r ^{2}=30/120=1/4\]\[r ^{2}=1/4\]

OpenStudy (whpalmer4):

\[120r^2=30\] I think you mean...

OpenStudy (radar):

Take it the next step.

OpenStudy (anonymous):

So its 30 instead of 130? Im confused.

OpenStudy (johnweldon1993):

the answer choice would be 60....but they are explaining exactly how to arrive at it :)

OpenStudy (anonymous):

Are you 100% sure?

OpenStudy (whpalmer4):

I was just correcting a typo in @radar's first line...was hoping to avoid you getting confused about where he got 130 from, but ended up confusing you about something else instead, can't win :-)

OpenStudy (johnweldon1993):

quite :) ...what radar and whpalmer were doing was showing you how the common ratio would be calculated as 1/2 ...and you would have used that to solve this :)

OpenStudy (anonymous):

Its 60, not 60(+-)?

OpenStudy (anonymous):

I need to know for sure.

OpenStudy (radar):